OpenStudy (anonymous):

Given one coordinate of a point on the unit circle, (√2\2, y), find the missing coordinate. A) 1\2 B) √2\2 C) 0 D) 3\4

OpenStudy (anonymous):

Ooh oh oh I just learned this one sec

OpenStudy (jdoe0001):

the UNIT circle means, the circle has a radius of 1 so \(\bf \left(\frac{\sqrt{2}}{2}, y\right)\to (x,y)\to \begin{cases} x=\frac{\sqrt{2}}{2} \\ \quad \\ y=? \end{cases}\)  thus